OpenStudy (anonymous):

qustion on ratio.... if a team consist of 20 members in which 20 % are mathematician then how many member should they hire so that the ratio would be 10/11

hartnn (hartnn):

so, before hiring, how many mathematicians are there out of 20 ?

OpenStudy (anonymous):

4

hartnn (hartnn):

goo, so old ratio = 4/20 new ratio = 10/11 what would you add to both numerator and denominator to change the ratio from old to new ?

hartnn (hartnn):

like solve \(\dfrac{4+x}{20+x}=\dfrac{10}{11}\)

OpenStudy (anonymous):

@hartnn: ok thnks i got it now... :)

hartnn (hartnn):

cool :) welcome ^_^